# Vision Foundation Models Can Be Good Tokenizers for Latent Diffusion Models

### v1

@@ 4.1 Setup
SD-VAE 60M 0.58 207.6 80 7.13 - 0.003 0.004 0.005 -
VA-VAE 160M 0.30 213.6 64 5.14 130.2 0.009 0.202 0.135 $33.2\%$
VFM-VAE 44M 0.52 208.0 64 3.80 152.8 0.111 0.188 0.191 $\mathbf{+1.6\%}$
Figure 4: CKNNA comparison across layers of the generative models. (a) Without explicit VFM alignment, the diffusion model combined with VFM-VAE consistently achieves higher layer-wise CKNNA than other tokenizer baselines. (b) With alignment enabled, the VFM-VAE system further improves, surpassing other diffusion-aligned baselines in average CKNNA.
@@ 4.3 Impact on Diffusion Models: Representation Diagnose
CKNNA might not be sufficient to evaluate latent representation of visual tokenizer. Though CKNNA (Huh et al., 2024) has been used to evaluate representations within diffusion models, its effectiveness in assessing visual tokenizers remains underexplored. As shown in Table 2, both VA-VAE and VFM-VAE achieve similar CKNNA scores (computed with DINOv2-Giant); however, when integrated into generative models, VFM-VAE outperforms VA-VAE by a significant margin. This discrepancy suggests that CKNNA alone may not fully capture the representational properties that make a tokenizer effective for generative modeling.
@@
SE-CKNNA of latent representation reflects the representation quality in LDMs. In Figure 4a, we compare layer-wise CKNNA within generative models trained without explicit alignment. We find that VFM-VAE consistently achieves higher layer-wise CKNNA and superior generative performance. Moreover, while the VA-VAE baseline exhibits consistently weaker layer-wise CKNNA compared to its alignment upper bound (computed between DINOv2-Large and DINOv2-Giant), VFM-VAE combined with LightningDiT reaches a peak CKNNA of 0.46—corresponding to 85% of the upper bound (computed between SigLIP2-Large and DINOv2-Giant)—indicating that it inherits VFM properties more effectively. For the direct generation metric, VFM-VAE + LightningDiT achieves an FID of 3.80 at 64 epochs, significantly outperforming the VA-VAE baseline’s FID of 5.14 at the same diffusion training stage.
@@
Joint tokenizer and diffusion alignment yields consistently high-quality representations. Building on the strong results achieved without explicit alignment, we further explore combining VFM-VAE with alignment methods in diffusion models. As shown in Figure 4, when paired with LightningDiT, shallow-layer alignment remains weak, improving only around layer 16. Ideally, the model should maintain features well aligned with VFM throughout its depth. This raises a natural question: can we explicitly enforce VFM alignment in the shallow layers to ensure consistency across all depths?
@@ Appendix A CKNNA Evaluation Details
Following Huh et al. (2024), we fix $k=10$ and apply outlier filtering with channel-wise normalization before computing CKNNA. For layer-wise analysis in generative models, we use spatial tokens only (discarding [CLS] tokens as in (Yu et al., 2024; Wu et al., 2025)) and compute the alignment after global pooling along the spatial dimension for both model and reference features.
@@ C.1 Not Only One VFM Choice
VA-VAE SigLIP2-Large 51M (40 epochs) 0.099 0.84 207.4 7.83 115.1
@@ Appendix D Hyperparameter And More Implementation Details
Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$

### v2

@@ 4.1 Setup
SD-VAE 60M 0.58 207.6 80 7.13 - 0.004 0.004 0.005 -
VA-VAE 160M 0.30 213.6 64 5.14 130.2 0.213 0.202 0.135 $33.2\%$
VFM-VAE 44M 0.52 208.0 64 3.80 152.8 0.164 0.188 0.191 $\mathbf{+1.6\%}$
Figure 4: CKNNA comparison across layers of the generative models. (a) Without explicit VFM alignment, the diffusion model combined with VFM-VAE achieves higher average layer-wise CKNNA than other tokenizer baselines. (b) With alignment enabled, the VFM-VAE system further improves, consistently surpassing other diffusion-aligned baselines in CKNNA.
@@ 4.3 Impact on Diffusion Models: Representation Diagnose
CKNNA might not be sufficient to evaluate latent representation of visual tokenizer. Though CKNNA (Huh et al., 2024) has been used to evaluate representations within diffusion models, its effectiveness in assessing visual tokenizers remains underexplored. As shown As shown in Table 2, VA-VAE achieves a clearly higher CKNNA score (computed with DINOv2-Giant) than VFM-VAE; however, when integrated into generative models, VFM-VAE outperforms VA-VAE by a significant margin. This discrepancy suggests that CKNNA alone may not fully capture the representational properties that make a tokenizer effective for generative modeling.
@@
SE-CKNNA of latent representation reflects the representation quality in LDMs. In Figure 4a, we compare layer-wise CKNNA within generative models trained without explicit alignment. We find that VFM-VAE consistently achieves higher layer-wise CKNNA and superior generative performance. Moreover, although VFM-VAE performs slightly lower than VA-VAE at the early stages, it surpasses VA-VAE by a large margin in the later stages, leading to a higher average CKNNA overall. Its peak CKNNA reaches 0.52, exceeding the 0.50 reference level (computed between SigLIP2-Large and DINOv2-Giant), indicating that it inherits VFM properties more effectively. For the direct generation metric, VFM-VAE + LightningDiT achieves an FID of 3.80 at 64 epochs, significantly outperforming the VA-VAE baseline’s FID of 5.14 at the same diffusion training stage.
@@
Joint tokenizer and diffusion alignment yields consistently high-quality representations. Building on the strong results achieved without explicit alignment, we further explore combining VFM-VAE with alignment methods in diffusion models. As shown in Figure 4, when paired with LightningDiT, shallow-layer alignment remains weak, with CKNNA showing little improvement in the early layers. The scores begin to rise rapidly only after the 12th layer, and by the 16th layer, the VFM-VAE system surpasses the VA-VAE baseline in alignment quality. Ideally, the model should maintain features well aligned with VFM throughout its depth. This raises a natural question: can we explicitly enforce VFM alignment in the shallow layers to ensure consistency across all depths?
@@ Appendix A CKNNA Evaluation Details
Following Huh et al. (2024), we fix $k=10$ and apply outlier filtering with channel-wise normalization before computing CKNNA. For layer-wise analysis in generative models, we use spatial tokens only (discarding [CLS] tokens as in REG (Wu et al., 2025)) and compute the alignment after global pooling along the spatial dimension for both model and reference features. All generative models are evaluated under a unified setting, where the input is the noised latent at diffusion timestep $t=0.5$, conditioned on the null label, consistent with the setup in REPA (Yu et al., 2024).
@@ C.1 Not Only One VFM Choice
VA-VAE SigLIP2-Large 51M (40 epochs) 0.102 0.84 207.4 7.83 115.1
@@ Appendix D Hyperparameter And More Implementation Details
Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $5\times 10^{-5}$ $5\times 10^{-5}$
Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $5\times 10^{-5}$ $5\times 10^{-5}$